OpenStudy (anonymous):

Solve for x: −2x + 5 < 7 x > −1 x < −1 x > −6 x < −6

jigglypuff314 (jigglypuff314):

solve it like you would an equation :) −2x + 5 < 7 add 2x to both sides +2x +2x 5 < 7 + 2x subtract 7 from both sides -7 -7 -2 < 2x then divide both sides by 2 /2 /2

OpenStudy (anonymous):

First, let's work on the left hand side of your inequality, the -2x + 5 This means, for instance, to see if it can be simplified at all. Multiply x and 2 Multiply x and 1 The x just gets copied along. The answer is x x 2*x evaluates to 2x Because of the minus sign 2x becomes - 2x The answer is -2x -2*x+5 evaluates to -2x+5 So, all-in-all, the left hand side of your inequality can be written as: -2x+5 Now, let's work on the right hand side of your inequality, the 7 The right hand side of your inequality can be written as: 7 So with these (any) simplifications, the inequality we'll set out to solve is: -2x+5 ‹ 7 Move the 5 to the right hand side by subtracting 5 from both sides, like this: From the left hand side: 5 - 5 = 0 The answer is -2x From the right hand side: 7 - 5 = 2 The answer is 2 Now, the inequality reads: -2x ‹ 2 To isolate the x, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the x on the left side of the inequality. The last step is to divide both sides of the inequality by -2 like this: To divide x by 1 The x just gets copied along in the numerator. The answer is x -2x ÷ -2 = x 2 ÷ -2 = -1 Note! Here's an important point about inequalities: We are now dividing by -2, which is a number less than zero! This changes the form of your inequality. Or, in other words, your ‹ will now become a ›. The solution to your inequality is: x › -1 So, your solution is: x must be greater than -1

OpenStudy (anonymous):

-1

OpenStudy (the_fizicx99):

There's no need to move them around that much. −2x + 5 < 7 -5 -5 -2x < 2 /-2 /-2 x > -1 When you divide or multiply by a negative you flip the sign.

OpenStudy (anonymous):

wow thanks guys. :)

OpenStudy (the_fizicx99):

@mangorox why do you keep copying and pasting?

jigglypuff314 (jigglypuff314):

in that case, you type really really fast 0.o

OpenStudy (anonymous):

lol you do

OpenStudy (anonymous):

Yeah...I type fast...also because I've done this question for homework before so...

OpenStudy (anonymous):

oh nice

OpenStudy (the_fizicx99):

1) Why did you delete the post you tagged me? 2) How do you type 250 words in like 2 seconds? o.o

OpenStudy (anonymous):

@tHe_FiZiCx99 1) I didn't delete the post, must be something with OpenStudy. 2) I was typing shortly after @jigglypuff314 was typing...so for about 1-2 minutes.

OpenStudy (the_fizicx99):

Lol what?! The question opened and jiggly puff was already typing ._.

OpenStudy (anonymous):

@tHe_FiZiCx99 I said I started typing after jiggly puff posted his comment...so I typed for more than "2 seconds". Probably about 1-2 minutes...

OpenStudy (the_fizicx99):

o.o ok...